OpenStudy (anonymous):

(7m+3n) (5m-6n) need help.

OpenStudy (karatechopper):

Do you know how to do the distributive property?

OpenStudy (anonymous):

yes.. but can you explain me what to do with he letters?

OpenStudy (karatechopper):

The letters will stick right with their coefficient

OpenStudy (karatechopper):

7m is all one term.

OpenStudy (anonymous):

hint pls,,

OpenStudy (karatechopper):

uhm, you will distribute 7m to the terms in the separate pair of parenthesis. Like 7m will distribute to 5m and to -6n

OpenStudy (karatechopper):

same goes for 3n

OpenStudy (anonymous):

draw it pls.. :(

OpenStudy (anonymous):

should i multiplied it?

hero (hero):

The distributive property is a(b + c) = ab + ac In this case, treat (7m + 3n) as "a" treat (5m - 6n) as "b + c" Then: (7m + 3n)(5m - 6n) = 5m(7m + 3n) - 6n(7m + 3n) <---Apply distributive property again = 35m^2 + 15mn - 42mn - 18n^2 = 35m^2 - 27mn - 18n^2

OpenStudy (anonymous):

hero pls..

OpenStudy (anonymous):

Thanks my answer is correct :)

OpenStudy (anonymous):

no its ok :)

OpenStudy (pinksapphire):

But yeah hero is right :)

OpenStudy (anonymous):

yeah :) Thanks to all of u ~

OpenStudy (pinksapphire):

Okay, sorry though for confusing you...today has been a long day for me...

OpenStudy (anonymous):

no.. its ok :)

OpenStudy (pinksapphire):

Alright :)